# The Cost of Dynamic Reasoning: Demystifying AI Agents and Test-Time Scaling from an AI Infrastructure Perspective

### v1

@@ IV Demystifying AI Agents
In this section, we demystify the execution characteristics of AI agents. Section IV-A first examines an agent’s single-request execution, analyzing how iterative LLM calls and interactions with external tools affect latency and resource usage. This is followed by a detailed exploration of the LLM inference and tool-calling characteristics of agents in Section IV-B. Lastly, Section IV-C shifts the focus to the serving environment of agentic systems where multiple requests are handled concurrently, identifying system-level bottlenecks and scalability issues that emerge during the deployment of agents.
@@ IV-A Overall Workflow of AI Agents
An AI agent’s workflow alternates between two main stages: LLM inference and tool use (Section II-B). This process forms a dynamic feedback loop in which the LLM determines the next action, invokes an external tool if necessary, and incorporates the resulting observation into the next inference step. This section analyzes the execution patterns of AI agent workflows, with a focus on their system-level characteristics.
@@
Effect of LLM and tool calls on latency. Figure 4 shows the average number of LLM and tool invocations per request across benchmarks. While CoT performs only a single LLM inference per request, tool-augmented agentic systems—ReAct, Reflexion, LATS, and LLMCompiler—require significantly more LLM calls, averaging 9.2 times more than CoT. Among these, LATS exhibits the highest LLM invocation count, with an average of 71.0 LLM calls per request. This is primarily due to its use of tree search, which explores multiple reasoning branches (i.e., child nodes) by issuing separate LLM inferences for each one when expanding a tree node.
@@
Figure 5 presents the end-to-end latency and the latency breakdown of each agent’s execution. While most agents exhibit a similar number of LLM and tool calls per request (Figure 4), the latency contribution from tool calls varies significantly depending on the workload. This discrepancy is primarily due to differences in the underlying tool execution latencies. For example, WebShop uses lightweight tools that interact with locally hosted webpages, resulting in tool latencies as low as 20 ms per call. In contrast, HotpotQA relies on the Wikipedia API, where individual calls take an average of 1.2 seconds. As a result, tool execution dominates the overall latency breakdown in this case. We discuss tool-calling characteristics further in Section IV-B.
@@
On average, LLM inference and tool execution account for 69.4% and 30.2% of total latency, respectively. Both stages contribute significantly to overall latency, but they are difficult to overlap due to their sequential dependency. Specifically, the LLM output is needed to determine which tool to call and with what parameters. Conversely, the next LLM invocation typically relies on the observation returned by the tool. Although LLMCompiler attempts to mitigate this dependency by streaming intermediate plans to the scheduler for asynchronous execution of tool calls (thus concurrently executing it with planning), the observed overlap accounts for only 18.2% of total latency.
Key takeaway #1: The sequential dependency between LLM inference and tool execution creates a fundamental bottleneck in an AI agent’s execution latency, as these stages cannot be easily parallelized. To improve the end-to-end performance of AI agents, future systems should explore techniques that reduce serialization between LLM inference and tool execution. For example, asynchronous pipelines or speculative tool invocation could help overlap LLM inference with tool execution—a capability partially achieved with LLMCompiler. Optimizing either the LLM or the tool in isolation is insufficient; meaningful end-to-end latency reduction requires coordinated, system-level optimization.
@@
Figure 6: Breakdown of GPU runtime by usage (left axis, bar graph) and the resulting average GPU utilization (right axis, diamond marker). GPU utilization is measured as the fraction of actively used GPU cores, using NVIDIA’s DCGM [27].
@@
Agentic workflow’s effect on GPU compute utility. Figure 6 breaks down GPU runtime by usage and shows the resulting average GPU utilization while handling a single request. Compared to CoT, which performs a single LLM inference without any external interaction, tool-augmented AI agents exhibit significantly lower GPU utilization. This is primarily due to the tool execution phases, which introduce GPU idle periods. With the exception of HumanEval’s test generation tool—which uses the GPU to execute an LLM—the tools in our experiments generally run entirely on the local CPU or external systems (e.g., accessing the Wikipedia webpage). As a result, the GPU typically remains idle during tool-calling periods, leading to as much as 54.5% of the execution time being idle. When the GPU is executing the LLM, its activity can be further divided into the prefill and decode stages [37, 67, 28], which account for 4.7% and 74.1% of the GPU’s execution time, respectively. As noted in  [1, 21, 6, 2], the decode stage is known to be memory-bound. Consequently, the large fraction of time spent in the decode stage further contributes to the underutilization of GPU resources.
@@
Because the sequential dependency between LLM inference and tool calls limits parallel execution opportunities within a single request (i.e., intra-request parallelism), improving overall resource utilization requires leveraging inter-request parallelism. We explore this direction in Section IV-C, where we discuss the implications of serving AI agents over multiple queries with LLM request batching [66, 19, 53].
Key takeaway #2: AI agents interleave GPU-bound LLM inference with CPU- or I/O-bound tool calls, resulting in non-trivial GPU idle time. This underutilization arises from the inherently sequential workflow of AI agents, which limits the potential for intra-request parallelism. To address this, agent-serving systems should adopt inter-request concurrency mechanisms to better utilize idle GPU cycles and improve overall throughput.
@@
Figure 7: Latency distribution when serving a non-agentic ShareGPT workload and a ReAct-based agent. Latency is measured while processing one request at a time, with prefix caching enabled.
@@
Latency distribution when serving AI agents. Figure 7 compares the end-to-end latency distributions of a conventional, non-agentic LLM service using ShareGPT and a ReAct-based agent system. The ShareGPT dataset represents a typical chatbot workload, where each response is generated by a single LLM inference. As shown, this results in a relatively low and consistent latency distribution, with most responses completing within 3 to 7 seconds. In contrast, the ReAct-based agent exhibits a much broader latency distribution with a heavier tail. This is due to its multi-step reasoning and reliance on external tool usage. Because the number of reasoning steps and tool calls varies across requests in AI agents, the associated computational demands also fluctuate. As a result, there is significant variance in latency across queries targeting AI agents.
Key takeaway #3: Unlike a single-pass LLM service with predictable latency, an agent’s iterative reasoning leads to highly variable, multi-step execution times, i.e., a heavy-tailed latency distribution per query. Systems must therefore be designed to handle this per-request variability. Techniques such as adaptive scheduling or elastic resource allocation can potentially help maintain efficiency and responsiveness, even as reasoning depth varies significantly across queries.
@@ IV-B LLM Inference and Tool-Calling Characteristics
Figure 8: Breakdown of input and output tokens in LLM inference. Instruction and Few-shot (light and dark gray) represent input tokens that are statically fixed as part of the initial prompt to the LLM. User (black) denotes input tokens provided by the user as part of the query. Output (red) refers to tokens generated by each LLM call. LLM history (green) and Tool history (yellow) represent tokens accumulated from previous LLM outputs and tool responses, respectively, which are then included as input tokens during the next LLM call.
Figure 9: Breakdown of token count per each AI agent’s iterative reasoning step for HotpotQA. While Instruction and Few-shot tokens remain constant, the accumulation of LLM history and Tool history tokens significantly increases the input context length over time.
@@
Compared to CoT, other AI agents consume more input tokens but generate fewer output tokens per LLM call. This is because these agents decompose their reasoning into multiple steps, invoking the LLM several times and incorporating tool usage between LLM calls. As a result, the total output is spread across multiple LLM invocations. In contrast, CoT completes its reasoning in a single LLM call, producing a longer, continuous output in just a single step. Among the AI agents, LATS stands out for generating notably longer output sequences. This is because it produces multiple output samples from a single input to generate candidate expansions when growing a tree node.
@@
Although the ratio of LLM and tool history tokens varies across workloads, most benchmarks exhibit substantial growth in input history over multiple iterations. An exception is LATS, which includes only the path from the root to the current node, rather than concatenating all prior interaction histories. As shown in Figure 9, initial inputs are typically around 1,000 tokens, but the input size increases to 3–4$\times$ as prior LLM outputs and tool responses are appended to the input context of subsequent LLM calls. Because histories accumulate sequentially, consecutive LLM calls share common prefixes in their input contexts. These long input contexts result in high KV cache usage per request and considerable prefix overlap across iterations. This behavior presents an opportunity to improve GPU compute and memory efficiency through prefix caching [19], as detailed below.
Key takeaway #4: AI agents accumulate long input contexts over iterative reasoning steps by appending prior LLM outputs and tool responses. This behavior results in large KV cache footprints and increasing GPU memory usage. However, because many of these tokens are shared across iterations, system-level optimizations such as prefix caching are needed to reduce redundant computation and help alleviate memory overhead in multi-step reasoning workloads.
Figure 10: Breakdown of LLM inference latency with and without prefix caching enabled.
Figure 11: LLM inference latency with and without prefix caching.
@@
Figure 10 shows the proportion of prefill and decoding latency during LLM inference, with and without prefix caching. For CoT, LLM inference occurs only once per request, and the shared prefix across inferences is minimal. Moreover, CoT typically generates a relatively large number of output tokens, making decoding the dominant contributor to latency. In contrast, AI agents operate iteratively and accumulate long input contexts due to interaction histories. As a result, prefix caching reduces prefill latency by an average of 58.6%, demonstrating its effectiveness in improving compute efficiency by avoiding redundant computations through prefix reuse.
As shown in Figure 11, the overall impact of prefix caching on end-to-end LLM inference latency varies significantly by workload type. In CoT-style prompting, the decoding phase dominates execution time—a characteristic common to virtually all static reasoning models (Figure 1(a,b))—leaving the prefill phase as only a small portion of the total. As a result, prefix caching offers limited latency reduction for CoT. In contrast, agentic workloads exhibit a non-negligible prefill phase (highlighted by the red bars in Figure 10) due to the accumulation of long input contexts over iterative steps. This makes them more amenable to performance gains from prefix caching, which eliminates redundant computation in the prefill stage and yields an average 15.7% reduction in end-to-end latency. While this per-request improvement may seem modest, the reduction in prefill time can significantly alleviate system-level bottlenecks. In token-level schedulers like vLLM, long prefill phases can delay the scheduling of concurrent requests. By shortening these phases, prefix caching can improve scheduling efficiency and increases overall system throughput. This effect is examined further in Section IV-C (Figure 15).
Key takeaway #5: Prefix caching reduces redundant prefill computation across iterative LLM calls, improving per-call compute efficiency and shortening scheduling-critical prefill phases. However, because decoding dominates overall latency and is less amenable to caching, the net reduction in end-to-end inference time remains modest. Therefore, system-level optimizations such as speculative token generation or leveraging reduced prefill time to increase batching opportunities are essential complements to caching for improving overall performance.
Figure 12: Average GPU memory requirement for KV cache per AI agent request, with and without prefix caching.
Effect of prefix caching on AI agent’s memory efficiency. While the previous analysis demonstrated how prefix caching improves compute efficiency by bypassing prefill operations for shared prefix tokens, we now examine its impact on GPU memory requirements by measuring the average GPU memory required to store the KV cache (Figure 12). On average, tool-augmented AI agents consume 3.0$\times$ more memory per request than CoT, and up to 5.4$\times$ more in the worst case—even with prefix caching. This overhead arises from the iterative nature of agent workflows, where each LLM call appends intermediate reasoning steps and tool responses to the context, resulting in an increasingly large input for each LLM inference.
@@
Key takeaway #6: Iterative reasoning in tool-augmented AI agents substantially increases GPU memory usage due to the accumulation of long input contexts across multiple LLM calls. This highlights the need for memory-optimized system designs that enable context reuse through shared KV cache and potentially leverage specialized hardware or heterogeneous memory hierarchy for efficient context storage and retrieval (e.g., offloading all or parts of KV cache contexts to CPU memory or SSD).
@@ IV-C AI Agent Serving Characteristics
Figure 13: High-level overview of our AI agent serving system.
@@
To examine the characteristics of AI agent serving, we implement an agent serving system, as illustrated in Figure 13. When a user sends a request to the agent server’s entry point, each worker processes the request according to the agent’s workflow. Depending on the current step of the task, a worker either sends a request to the LLM inference server or executes a tool. Tool execution may occur locally (e.g., code interpreters, custom functions) or involve external resources (e.g., web search, API calls). Each worker operates asynchronously, and LLM inference requests from multiple workers can be batched at the LLM backend (e.g., vLLM) for high-throughput processing using continuous batching [66, 19]. We adopt vLLM’s default first-come-first-served (FCFS) scheduler in the LLM inference backend. To simulate realistic traffic, input queries to the agent server are randomly sampled and issued to the server following a Poisson arrival distribution [26].
Figure 14: 50th and 95th percentile latencies for chatbot (ShareGPT) and AI agent (ReAct) workloads as queries per second (QPS) rates increase, with prefix caching enabled.
Importance of concurrent request scheduling. Before comparing AI agent serving against conventional static reasoning based chatbot (ShareGPT) serving scenarios, we first highlight the importance of concurrently servicing AI agent requests. When ReAct agents are executed sequentially, the average latency is 9.6 seconds for HotpotQA and 5.3 seconds for WebShop, limiting throughput to 0.10 and 0.19 QPS, respectively. With concurrent execution, throughput improves to 2.6 and 1.2 QPS for HotpotQA and WebShop, respectively, achieving 25$\times$ and 6.2$\times$ gains at the cost of a 2.1$\times$ increase in average latency. The greater throughput gain in HotpotQA arises from its longer tool latency, which causes the GPU to remain idle for extended periods. These idle intervals can be effectively utilized by executing other requests, enabling higher concurrency and improved throughput.
Comparison with conventional static reasoning LLM services. We now compare AI agent serving with a conventional LLM serving scenario, represented by the chatbot (ShareGPT) workload. ShareGPT, a typical single-turn LLM service, processes user queries in a single inference pass. Figure 14 shows the changes in end-to-end tail latencies for chatbot (ShareGPT) and AI agent (ReAct) workloads as input queries per second (QPS) to the server increase. The peak throughput is measured as the maximum sustainable QPS at the knee of the tail latency curve. As depicted, the peak throughput of ReAct is significantly lower than that of ShareGPT. While ShareGPT can sustain up to 6.4 QPS, ReAct supports only 2.6 QPS on HotpotQA and 1.2 QPS on WebShop. This limitation stems from ReAct’s multi-step reasoning, where each request involves multiple LLM calls and tool interactions, significantly increasing latency.
Another point worth noting is that ReAct’s 95th percentile latency increases much more sharply with load. For every 1 QPS increase from 50% to 100% of peak throughput, the 95th percentile latency rises by 17.8 seconds on HotpotQA and 6.1 seconds on WebShop, whereas ShareGPT’s latency increases by only 0.9 seconds. This steep increase in tail latency highlights AI agent’s high sensitivity to load and the difficulty of maintaining low tail latency under heavy traffic. Overall, these observations underscores the need for careful scheduling and resource management when serving AI agents, especially compared to conventional LLM services.
Key takeaway #7: Compared to conventional LLM services, AI agent serving experiences lower throughput and significantly higher sensitivity to tail latency due to their dynamic execution patterns and iterative reasoning. This heightened sensitivity underscores the need for agent-specific scheduling and resource management strategies, such as agent-aware request dispatching or adaptive compute resource allocation per request, to maintain quality of service (QoS).
@@
Figure 15: 95th percentile latency for chatbot (ShareGPT) and AI agent (ReAct) workloads as QPS rates increase, with (solid line) and without (dashed line) prefix caching enabled.
@@
Figure 16: (a) Average and (b) maximum memory used in allocating KV caches, with and without prefix caching enabled. Evaluation is conducted at 0.2 QPS (HotpotQA) and 0.1 QPS (WebShop) data points using ReAct.
Effect of prefix caching on AI agent serving throughput. Prefix caching is an important system-level optimization that reduces redundant computation during the prefill phase of LLM inference by reusing previously computed key-value (KV) caches. While its impact on the latency of individual LLM calls is modest, it can substantially improve throughput and overall serving efficiency, particularly in AI agent serving.
Figure 15 compares the effect of prefix caching on chatbot (ShareGPT) and agentic (ReAct) workloads. ShareGPT shows only a modest 1.03$\times$ throughput improvement, as it performs a single LLM call per request with minimal repetition. In contrast, ReAct benefits significantly, achieving an average 5.62$\times$ increase in throughput. This is because agent workloads involve multiple LLM calls per request, amplifying the benefits of avoiding redundant prefill operations.
Key takeaway #8: Prefix caching significantly improves serving efficiency for AI agents by reducing prefill overhead and inter-request interference, resulting in much greater throughput gains than in conventional LLM workloads. Therefore, incorporating sophisticated caching mechanisms such as solutions that persist and reuse prefixes across queries is a promising direction for sustaining high throughput in complex, iterative reasoning tasks.
Effect of prefix caching on AI agent’s memory usage. We now investigate the impact of prefix caching on GPU memory efficiency in AI agent serving, focusing specifically on its effect on key-value (KV) cache size, one of the most significant contributors to memory usage in LLM inference. Figure 16 shows the GPU memory consumed in allocating KV caches, with and without prefix caching enabled, under identical QPS conditions. With prefix caching enabled, the average and maximum KV cache memory usage decrease by 51.7% and 63.5%, respectively, indicating improved memory efficiency. This reduction arises from the ability of prefix caching to reuse key-value pairs of shared prefix tokens across multiple LLM invocations across AI agent requests. Thus, prefix caching not only improves compute efficiency by eliminating redundant prefill operations but also reduces the KV cache memory footprint, enabling more efficient utilization of GPU memory during AI agent serving.
Figure 17: 95th percentile latency (left axis, solid lines) and prefix cache hit rate (right axis, dashed lines) when the GPU memory reserved for KV cache allocation is changed. Legends denote the GPU memory reserved for KV cache allocation, the value of which denotes the reserved memory size relative to the LLM model weight size.
Next, we analyze how the effectiveness of prefix caching and the serving performance of AI agents are impacted by the GPU memory capacity available for KV cache allocation. Figure 17 illustrates how varying the amount of GPU memory reserved for the key-value (KV) cache affects serving throughput. As shown in the figure, reducing the KV cache allocation pool significantly limits throughput, as the system cannot accommodate many concurrent requests. Compared to the 200% configuration, the maximum sustainable throughput drops by 86.3% under the 10% configuration and by 73.6% under the 20% configuration. Because the opportunity to batch multiple requests is limited under these two configurations, both exhibit noticeable queuing delays, as request scheduling becomes increasingly serialized. Under the 30% configuration, the KV cache allocation pool is just large enough to enable batching and minimize queuing delays; however, concurrent requests still compete for limited KV cache. This results in cache thrashing, which reduces the effectiveness of prefix reuse. As a result, the 30% configuration achieves 35% and 18% lower throughput for HotpotQA and WebShop, respectively, compared to the 200% configuration. This demonstrates that even moderate constraints in GPU memory can critically limit the serving performance of AI agents.
Key takeaway #9: Prefix caching improves GPU memory efficiency in AI agent serving by reusing shared prefixes across requests. However, due to the long and growing contexts in agent workloads, limited KV cache capacity can lead to cache thrashing and sharply degrade throughput, even under moderate memory constraints. This highlights the need for more intelligent cache management strategies, such as locality-aware hierarchical caching or KV cache compression techniques, that helps prevent cache saturation and sustain performance under concurrent agent requests.

### v2

@@ IV Demystifying AI Agents
Section IV-A first examines an agent’s single-request execution, followed by a detailed exploration of the LLM inference and tool-calling characteristics of agents in Section IV-B. Lastly, Section IV-C shifts the focus to the serving environment of agentic systems where multiple requests are handled concurrently, identifying system-level bottlenecks and scalability issues that emerge in agent deployment.
@@ IV-A Overall Workflow of AI Agents
Effect of LLM and tool calls on latency. Figure 4 shows the average number of LLM and tool invocations per request across benchmarks. While CoT performs only a single LLM inference per request, tool-augmented agentic systems require significantly more LLM calls, averaging 9.2 times more than CoT. Among these, LATS exhibits the highest LLM invocation count, with an average of 71.0 LLM calls per request. This is primarily due to its use of tree search, which explores multiple reasoning branches (i.e., child nodes) by issuing separate LLM inferences for each one when expanding a tree node.
@@
Figure 5 presents the end-to-end latency and the latency breakdown of each agent’s execution. While most agents exhibit a similar number of LLM and tool calls per request (Figure 4), the latency contribution from tool calls varies significantly depending on the workload. This discrepancy is primarily due to differences in the underlying tool execution latencies. For example, WebShop uses lightweight tools that interact with locally hosted webpages, resulting in tool latencies as low as 20 ms per call. In contrast, HotpotQA relies on the Wikipedia API, where individual calls take an average of 1.2 seconds. As a result, tool execution dominates the overall latency breakdown in this case.
@@
On average, LLM inference and tool execution account for 69.4% and 30.2% of total latency, respectively. Both stages contribute significantly to overall latency, but they are difficult to overlap due to their sequential dependency. Specifically, the LLM output is needed to determine which tool to invoke along with the required arguments. Conversely, the next LLM invocation typically relies on the observation returned by the tool. Although LLMCompiler attempts to mitigate this dependency by streaming intermediate plans to the scheduler for asynchronous execution of tool calls (thus concurrently executing it with planning), the observed overlap accounts for only 18.2% of total latency.
@@
Figure 6: Breakdown of GPU runtime by usage (left axis, bar graph) and the resulting average GPU utilization (right axis, diamond marker). GPU utilization is measured as the fraction of actively used GPU cores, using NVIDIA’s DCGM [dcgm].
@@
Agentic workflow’s effect on GPU compute utility. Figure 6 breaks down the GPU runtime by usage and reports the resulting average GPU utilization when handling a single request. Although this setup assumes the processing of a single agent task, concurrent LLM calls can be opportunistically batched to improve GPU utilization, whenever possible, to more efficiently execute agents such as LATS. Unlike CoT, which performs a single LLM inference without external interaction, it is possible for agents to experience longer GPU idle periods due to tool execution. The duration of these idle periods depends on the tool’s latency and whether it leverages the GPU. In WebShop, the tool interacts with locally hosted synthetic web pages, resulting in very short tool latencies (Figure 5), so agents do not experience notably higher GPU idle time (i.e., lower GPU utilization). HumanEval exhibits longer tool execution times (Figure 5), but the proportion of GPU idle time remains minimal because the tool it calls (which is the test generation tool) utilizes the GPU for LLM execution. In contrast, HotpotQA and MATH employ tools that operate on local CPUs or external systems, leading to substantial GPU idle periods that account for up to 54.5% of total execution time, resulting in significantly lower GPU utilization compared to CoT. When the GPU is executing the LLM, its activity can be further divided into the prefill and decode stages, which account for 4.7% and 74.1% of the GPU’s execution time, respectively. As noted in  [sarathi, Leviathan2023SpecDecode, medusa, ainslie2023gqa], the decode stage is known to be memory-bound. Consequently, the large fraction of time spent in the decode stage further contributes to the underutilization of GPU resources.
@@
Because the sequential dependency between LLM inference and tool calls limits parallel execution opportunities within a single request (i.e., intra-request parallelism), improving overall resource utilization requires leveraging inter-request parallelism. We explore this direction in Section IV-C, where we discuss the implications of serving AI agents over multiple queries with LLM request batching [orca, vllm, vllm_repo].
@@
Figure 7: End-to-end latency distribution of a non-agentic ShareGPT workload and ReAct-based agents with HotpotQA and WebShop workloads. Latency is measured while processing one request at a time, with prefix caching enabled.
@@
End-to-end latency distribution of AI agents. Figure 7 compares the end-to-end latency distributions of a conventional, non-agentic LLM service using ShareGPT and a ReAct-based agent system using HotpotQA and WebShop. The ShareGPT dataset represents a typical chatbot workload, where each response is generated by a single LLM inference. As shown, this results in a relatively low and consistent latency distribution, with most responses completing within 9.7 seconds. In contrast, the ReAct-based agent exhibits a much broader latency distribution with a heavier tail, due to its multi-step reasoning and reliance on external tools. Because the number of reasoning steps and tool calls varies across requests in agents, the associated computational demands also fluctuate. Consequently, there is significant variance in latency across queries targeting agents.
@@ IV-B LLM Inference and Tool-Calling Characteristics
Figure 8: Breakdown of input and output tokens in LLM inference. Instruction and Few-shot (light and dark gray) represent input tokens that are statically fixed as part of the initial prompt to the LLM. User (black) denotes input tokens provided by the user as part of the query. LLM history (green) and Tool history (yellow) represent tokens accumulated from previous LLM outputs and tool responses, respectively, which are then included as input tokens during the next LLM call. Output (red) refers to tokens generated by each LLM call.
Compared to CoT, AI agents generally have longer input tokens. This is because their inputs include additional elements such as agent role-aligned instructions (e.g., LLMCompiler requires instructions to generate a structured plan) and the accumulated context of previous LLM and tool interactions. For output tokens, each LLM call in agent workflows often generates fewer tokens than CoT, except for LATS. This is because agents typically decompose a single task into multiple steps, distributing the overall output across several LLM calls. In contrast, LATS often generates much longer outputs than CoT due to its workflow, where a single LLM call produces multiple candidate samples to expand the tree node.
@@
[…]
@@
Figure 9: Inference latency with(out) prefix caching and its breakdown.
Figure 9 shows LLM inference latency and its proportion of prefill and decoding latency, with and without prefix caching. For CoT, LLM inference occurs only once per request, and the shared prefix across inferences is minimal. Moreover, CoT typically generates a relatively large number of output tokens, making decoding the dominant contributor to latency. In contrast, AI agents operate iteratively and accumulate long input contexts due to interaction histories. As a result, prefix caching reduces prefill latency by an average of 60.1%, demonstrating its effectiveness in improving compute efficiency by avoiding redundant computations through prefix reuse. Beyond the prefill phase, prefix caching can also indirectly improve decoding efficiency. In systems that execute multiple parallel LLM calls, decoding may be stalled by ongoing prefill operations. By reducing prefill latency, prefix caching shortens this blocking period, thereby enabling faster decoding and explaining the larger speedup observed in LATS.
Overall, the impact of prefix caching on end-to-end LLM inference latency varies by workload type. While CoT workloads benefit less due to their decoding-dominant property, agentic workloads experience an average 15.7% reduction in end-to-end latency due to the accumulation of long input contexts over iterative steps. While this per-request improvement may seem modest, the reduction in prefill time can significantly alleviate system-level bottlenecks. In token-level schedulers like vLLM, long prefill phases can delay the scheduling of concurrent requests. By shortening these phases, prefix caching can improve scheduling efficiency and increase overall system throughput. This effect is examined further in Section IV-C (Figure 11).
[…]
@@
Figure 10: High-level overview of our AI agent serving system.
@@ IV-C AI Agent Serving Characteristics
To examine the characteristics of AI agent serving, we implement an agent serving system, as illustrated in Figure 10. When a user sends a request to the agent server’s entry point, each worker processes the request according to the agent’s workflow. Depending on the current step of the task, a worker either sends a request to the LLM inference server or executes a tool. Tool execution may occur locally (e.g., code interpreters, custom functions) or involve external resources (e.g., web search, API calls). Each worker operates asynchronously, and LLM inference requests from multiple workers can be batched at the LLM backend (e.g., vLLM) for high-throughput processing using continuous batching [orca, vllm]. We adopt vLLM’s default first-come-first-served (FCFS) scheduler in the LLM inference backend. To simulate realistic traffic, input queries to the agent server are randomly sampled and issued to the server following a Poisson arrival distribution [mlperf_datacenter].
[…]
@@
Figure 11: 95th percentile latency for chatbot (ShareGPT) and ReAct-based AI agents (HotpotQA and WebShop) as QPS rates increase, with (solid line) and without (dashed line) prefix caching enabled.
Comparison with conventional static reasoning LLM services. We now compare an AI agent serving with a conventional LLM serving scenario, represented by the chatbot (ShareGPT) workload. ShareGPT, a typical single-turn LLM service, processes user queries in a single inference pass. Figure 11 shows the changes in end-to-end tail latencies for chatbot (ShareGPT) and ReAct-based AI agent (HotpotQA and WebShop) workloads as input QPS to the server increases. The peak throughput is measured as the maximum sustainable QPS at the knee of the tail latency curve. As depicted, the peak throughput of ReAct is significantly lower than that of ShareGPT. While ShareGPT can sustain up to 6.4 QPS, ReAct supports only 2.6 QPS on HotpotQA and 1.2 QPS on WebShop, even with prefix caching enabled. This limitation stems from ReAct’s multi-step reasoning, where each request involves multiple LLM calls and tool interactions, significantly increasing latency.
Effect of prefix caching on AI agent serving throughput. Prefix caching is an important system-level optimization that reduces redundant computation during the prefill phase of LLM inference by reusing previously computed key-value (KV) caches. While its impact on the latency of individual LLM calls is modest (Figure 9), it can substantially improve throughput and serving efficiency for AI agents.
[…]
@@
Figure 12: (a) Average and (b) maximum memory used for KV caches, with and without prefix caching. Evaluation is conducted at 0.2 QPS (HotpotQA) and 0.1 QPS (WebShop) data points using ReAct.
Effect of prefix caching on AI agent’s memory usage. We now investigate the impact of prefix caching on GPU memory efficiency in AI agent serving, focusing specifically on its effect on key-value (KV) cache size, one of the most significant contributors to memory usage in LLM inference. Figure 12 shows the GPU memory consumption for KV cache allocation, with and without prefix caching enabled, under identical QPS conditions. With prefix caching enabled, the average and maximum KV cache memory usage decrease by 51.7% and 63.5%, respectively, indicating improved memory efficiency. This reduction arises from the ability of prefix caching to reuse key-value pairs of shared prefix tokens across multiple LLM invocations across AI agent requests. Thus, prefix caching not only improves compute efficiency by eliminating redundant prefill operations but also reduces the KV cache memory footprint, enabling more efficient utilization of GPU memory during AI agent serving.